OpenStudy (desiraem):

3. Two boys want to use a seesaw and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

how are you going to go about solving for d2?

OpenStudy (desiraem):

this is the correct equation. w^1*d^1= w^2*d^2 where w^1 = the weight of the first boy, d^1 = distance of the first boy from the fulcrum. w^2= weight of the second boy. d^2= distance of the second boy from the fulcrum rewrite the formula to solve for d^2.

OpenStudy (anonymous):

\[wd=w^2d^2\] or \[w _{1}d _{1}=w _{2}d _{2}\]

OpenStudy (desiraem):

the bottom one!

OpenStudy (anonymous):

ok. so solve for d2 by dividing both sides by w2

OpenStudy (anonymous):

\[\frac{ w _{1}d _{1} }{ w _{2} }=\frac{ w _{2}d _{2} }{ w _{2} }\]

OpenStudy (anonymous):

the w2's cancel on the right side now and you get\[d _{2}=\frac{ d _{1}w _{1} }{ w _{2} }\]

OpenStudy (anonymous):

gotta run, hope that helped